OpenStudy (anonymous):

lim x->-2 of (x^3+8)/(x+2)

OpenStudy (anonymous):

The top part can be written as: (x+2)(x^2-2x+4) Then we see that the (x+2) terms cancel out leaving you with: (x^2-2x+4). Now i am sure you can compute that limit.

OpenStudy (anonymous):

not cancel out, i mean divide out

OpenStudy (anonymous):

we can also use L'Hôpital's rule and differentiate both statements and we get 3x^2/1=3x^2=12

OpenStudy (jamesj):

Why use a gun when you can use a finger?

OpenStudy (anonymous):

what?

OpenStudy (jamesj):

I'm saying that using l'Hopital's rule is a bit over the top here.

OpenStudy (anonymous):

okay then why would you mention it

OpenStudy (jamesj):

@Lagrange, I don't think you're following me. I'm saying your method is better here than Noureldin's proposition.

OpenStudy (zarkon):

I love using a sledgehammer to pound in a nail. :)

OpenStudy (anonymous):

as in "why use an elephant gun to kill and ant" point is that since you have a rational function and replacing x by -2 gives 0/0 tells you that it MUST factor as \[\frac{(x+2)\times \text{something}}{x+2}\]

OpenStudy (anonymous):

in other words if you want \[\lim_{x\rightarrow a}\frac{P(x)}{Q(x)}\] for polynomials P and Q and you find that \[P(a)=Q(a)=0\] the both numerator and denominator MUST factor as \[(x-a)\times \text{stuff}\] but the factor theorem.

OpenStudy (anonymous):

*by